# Hardness of Regular Expression Matching with Extensions

### v1

@@
The regular expression matching problem asks whether a given regular expression of length $m$ matches a given string of length $n$. As is well known, the problem can be solved in $O(nm)$ time using Thompson’s algorithm. Moreover, recent studies have shown that the matching problem for regular expressions extended with a practical extension called lookaround can be solved in the same time complexity. In this work, we consider three well-known extensions to regular expressions called backreference, intersection and complement, and we show that, unlike in the case of lookaround, the matching problem for regular expressions extended with any of the three (for backreference, even when restricted to one capturing group) cannot be solved in $O(n^{2-\varepsilon}\operatorname{poly}(m))$ time for any constant $\varepsilon>0$ under the Orthogonal Vectors Conjecture. Moreover, we study the matching problem for regular expressions extended with complement in more detail, which is also known as extended regular expression (ERE) matching. We show that there is no ERE matching algorithm that runs in $O(n^{\omega-\varepsilon}\operatorname{poly}(m))$ time ($2\leq\omega<2.3716$ is the exponent of square matrix multiplication) for any constant $\varepsilon>0$ under the $k$-Clique Hypothesis, and there is no combinatorial ERE matching algorithm that runs in $O(n^{3-\varepsilon}\operatorname{poly}(m))$ time for any constant $\varepsilon>0$ under the Combinatorial $k$-Clique Hypothesis. This shows that the $O(n^{3}m)$-time algorithm introduced by Hopcroft and Ullman in 1979 and recently improved by Bille et al. to run in $O(n^{\omega}m)$ time using fast matrix multiplication was already optimal in a sense, and sheds light on why the theoretical computer science community has struggled to improve the time complexity of ERE matching with respect to $n$ and $m$ for more than 45 years.
@@ 1 Introduction
On the other hand, regular expression matching becomes much harder when extensions are present. A prominent example is backreference, an extension that allows reusing previously matched substrings. The extension is practically popular and supported in the standard libraries of many popular programming languages including Java, Python, JavaScript and more. Unfortunately, we cannot solve the matching problem for regular expressions extended with backreference (rewbs for short) using Thompson’s algorithm because rewbs are strictly more expressive than regular languages (in fact, they are not even context-free) [10] and therefore an equivalent NFA does not in general exist. Instead, it can be solved by simulating memory automata, which are a model equivalent to rewbs introduced by Schmid [39]. Nevertheless, the resulting matching algorithm for rewbs with at most $\ell$ capturing groups takes $O(n^{2\ell+1}m)$ time [39, 40, 33], and at present we have an additional $n^{2}$ factor even when $\ell=1$.
@@
The other well-known extensions that make the matching problem difficult are intersection and complement, where regular expressions extended with each of them are called semi-extended regular expressions (semi-EREs) and extended regular expressions (EREs), respectively. We note that EREs essentially subsume semi-EREs because intersection can be expressed using complement and alternation with only a linear increase in the length of the expression. The best known time complexity of both semi-ERE matching and ERE matching with respect to $n$ and $m$ is $O(n^{\omega}m)$ [14], where $2\leq\omega<2.3716$ [44] is the exponent of square matrix multiplication.
@@
In light of the above, we pose the following question: Do these extensions actually increase the time complexity of regular expression matching, or do they only appear hard because we have not yet discovered the right algorithmic ideas? Recent studies [28, 21, 9] have shown a surprising fact that lookaround, which, like backreference, is a practically popular extension that tests surrounding contexts of a position and is often used to mimic intersection and complement, does not actually increase the complexity of the matching problem by proposing algorithms that solve the problem in $O(nm)$ time.
1.1 Result I: Hardness of Matching with Backreference and Intersection
@@
In this paper, we consider the three extensions described above, backreference, intersection and complement, and show that regular expression matching extended with each of the three (for backreference, even when restricted to one capturing group) is hard under the Orthogonal Vectors Conjecture (OVC). We note that this also holds under the Strong Exponential Time Hypothesis (SETH) [25] because SETH implies OVC [45, 47].
@@
Unless OVC fails, there is no algorithm that solves rewb matching in $O(n^{2-\varepsilon}\operatorname{poly}(m))$ time for any constant $\varepsilon>0$, even when restricted to one capturing group.
@@
Theorems 1 and 2 state a much stronger statement than merely that we should not hope for a matching algorithm for rewbs and semi-EREs that runs in $O(nm)$ time. Indeed, it states that if we want to solve rewb matching or semi-ERE matching in truly subquadratic time in $n$ (i.e., $n^{2-\varepsilon}$ for some $\varepsilon>0$), then we are forced to admit a superpolynomial blowup in $m$ unless OVC fails.
The results are of interest not only to theorists but also to practitioners. They provide evidence that the real-world regular expression engines which guarantee worst-case $O(nm)$-time complexity (e.g., that in the de facto standard regular-expression library for Rust [^1] There, $m$ denotes the length of the input regular expression after an extension called counted repetition has been eliminated; see “Untrusted input” in [1].) are all forced to give up the full support for backreference and intersection, and would need to somehow restrict their usage such that our reductions in the proofs of Theorems 1 and 2 cannot apply. Thus, the results are of great practical value because it provides a guideline on how far the expressiveness can be increased in regular expression engines while retaining the good complexity.
@@
1.2 Result II: Hardness of Matching with Complement
@@ 1.3 Technical Overview
All lower bounds shown in this paper are of the form $O(n^{c-\varepsilon}\operatorname{poly}(m))$ and therefore do not rule out algorithms that run in superpolynomial time in $m$ (e.g., $O(n^{c-1}2^{m})$). Still, this polynomial-in-$m$ lower bound rules out algorithms that are efficient with respect to both $n$ and $m$. Moreover, restricting the time complexity of ERE matching with respect to $m$ is even necessary for deriving a meaningful lower bound on the complexity with respect to $n$ and $m$ because we could improve the complexity to be linear in $n$ if we were to allow the complexity to be unbounded in $m$. Indeed, this can be done by converting a given ERE to an equivalent NFA and running Thompson’s algorithm, at the cost of making the complexity nonelementary in $m$ [41]. Rosu has proposed a less extreme algorithm that runs in $O(n^{2}\,2^{m})$ time [37].
We prove our theorems by reductions from the respective problems. A key common technique in our reductions is to exploit the polynomial-in-$m$ lower-bound property to allow the regular expressions to be of polylogarithmic lengths with respect to the sizes of the instances to be reduced. For example, in the proofs of Theorems 1 and 2, we show reductions from the Orthogonal Vectors problem, which asks whether given sets $A$ and $B$ containing the same number of bit-vectors of dimension $d$ have a pair of orthogonal vectors $a\in A$ and $b\in B$. We encode the vectors in $A$ and $B$ into a string and construct a regular expression $r$ that nondeterministically guesses two vectors $a\in A$ and $b\in B$ from the string and checks their orthogonality. In proving a lower bound under OVC, we may assume that $d$ is logarithmic in the number of vectors in $A$ and $B$. Therefore, we can let the length of $r$ be polynomial in $d$ for proving our polynomial-in-$m$ lower bounds. As we shall show, such $r$ can be constructed by exploiting this property and the respective regular-expression extensions. A similar argument applies to Theorem 3 where we show a reduction from the $3k$-Clique problem.
@@ 2 Related Work
We first discuss the prior work on regular expression matching without extensions. A line of work has shown that the complexity of Thompson’s algorithm can be improved [29, 11, 12, 15, 16, 13]. On the other hand, Backurs and Indyk showed that the problem cannot be solved in $O((nm)^{1-\varepsilon})$ time for any $\varepsilon>0$ unless SETH fails [8]. Later, Abboud and Bringmann introduced a hypothesis, called the Formula-SAT Hypothesis, and showed a more fine-grained lower bound assuming it [3].
@@
Next, we discuss the prior work on rewb matching. Aho showed that the problem is NP-complete [6]. Moreover, it is known that rewbs can be regarded as a generalization of Angluin’s pattern languages (also known as patterns with variables), whose matching problem is already NP-complete [7, 18], even under bounded parameterizations [38, 19]. At this point, one may wonder about the novelty of Theorem 1 because NP-hardness already rules out rewb matching algorithms that run in $O(n^{2-\varepsilon}\operatorname{poly}(m))$ time (in fact, even those that run in time polynomial in $n+m$) assuming $\mathrm{P}\neq\mathrm{NP}$. However, it cannot rule out matching algorithms for rewbs with at most $\ell$ capturing groups for a fixed $\ell$, because the problem is solvable in $O(n^{2\ell+1}m)$ time as stated in the introduction and hence is in $\mathrm{P}$ when $\ell$ is fixed. By contrast, our Theorem 1 can rule out those for rewbs with even one capturing group under OVC. Furthermore, to our knowledge, there is no known implication between OVC and $\mathrm{P}\neq\mathrm{NP}$ and therefore these hardness results appear to be independent. Regarding upper bounds for the problem, Freydenberger and Schmid introduced a deterministic variant of rewbs together with algorithms that check whether a given rewb is deterministic in their sense in $O(m^{2})$ time and solve the matching problem for those rewbs in $O(m^{2}+n\ell)$ time [20]. Combining this with our Theorem 1, we obtain evidence under OVC that their determinism makes rewb matching strictly more tractable even when $\ell$ is fixed. Terauchi [42] has shown that the matching problem for rewbs satisfying a condition called constant degree ambiguity (CDA) can be solved in linear time in $n$. They leave open the decidability of checking whether a given rewb satisfies CDA. Additionally, Nogami and Terauchi [33] recently proposed an $O(n^{2}m^{2})$-time algorithm for a certain subclass of rewbs with one capturing group, improving the previously known best $O(n^{3}m)$ algorithm for the subclass with respect to $n$. We comment on the expressive power of rewbs. As mentioned in the introduction, backreference differs from intersection and complement in that it strictly increases the expressive power of regular expressions. The expressive power of rewbs is known to be incomparable with that of context-free languages (CFLs) [17, 10]. Furthermore, recent papers have shown that it is properly contained in that of EDT0L languages and incomparable with both the class of stack languages and that of multiple context-free languages (MCFLs), while a syntactic condition of rewbs, called the closed-star condition, decreases its expressive power to be properly contained in subclasses of those, namely the class of nonerasing stack languages and that of unary-MCFLs [31, 32, 34].
@@
Next, we discuss the prior work on ERE matching. As mentioned in the introduction, Hopcroft and Ullman [23] proposed in 1979 an algorithm that runs in $O(n^{3}m)$ time and no polynomial improvement over this with respect to $n$ and $m$ had existed for many years. ERE matching is extensively studied and has an interesting history. In fact, many papers had claimed to propose an improved $O(n^{2}m)$-time algorithm [22, 49, 27, 24] ([49] considers semi-ERE matching). However, all of these claims were later found to be incorrect; see [36, 37] for more details. Additionally, a number of studies had improved the complexity of ERE matching with respect to more refined parameters than just $n$ and $m$ [50, 51, 14]. As also mentioned in the introduction, Bille et al. [14] recently adopted fast matrix multiplication to improve the time and space complexity of Hopcroft and Ullman’s algorithm and that of the algorithm by Yamamoto and Miyazaki [51]. Furthermore, as mentioned in Section 1.3, Rosu proposed an algorithm that runs in time quadratic in $n$, but exponential in $m$ [37]. On the hardness side, Petersen proved that semi-ERE matching is LOGCFL-complete by showing a reduction $f$ from every context-free language (CFL) to semi-ERE matching [35]. Combining $f$, which is combinatorial, and the combinatorial reduction from $3k$-Clique to CFL recognition shown by Abboud et al. [2], we can derive the following hardness of semi-ERE matching: (i) there is no semi-ERE matching algorithm that runs in $O((n+m)^{\omega-\varepsilon})$ time for any $\varepsilon>0$ under the $k$-Clique Hypothesis, and (ii) there is no combinatorial semi-ERE matching algorithm that runs in $O((n+m)^{3-\varepsilon})$ time for any $\varepsilon>0$ under the Combinatorial $k$-Clique Hypothesis. However, note that these bounds cannot establish the optimality of Hopcroft and Ullman’s algorithm in the sense described in Section 1.2. By contrast, our Theorem 3 can do it. This is because our reduction outputs EREs of length only polylogarithmic in the input size, whereas $f$ outputs EREs of length linear in the input size.
@@
Finally, we note that our form of multivariate lower bounds is not new. For example, Abboud et al. proved that under OVC, for any $\varepsilon>0$, there is no $n^{2-\varepsilon}\,2^{o(k)}$-time algorithm that decides if the diameter of a given graph of treewidth $k$ is less than or equal to $2$ [5].

### v2

@@
The regular expression matching problem asks whether a given regular expression of length $m$ matches a given string of length $n$. As is well known, the problem can be solved in $O(nm)$ time using Thompson’s algorithm. Moreover, recent studies have shown that the matching problem for regular expressions extended with a practical extension called lookaround can be solved in the same time complexity. In this work, we consider four well-known extensions to regular expressions called backreference, squaring, intersection and complement, and we show that, unlike in the case of lookaround, the matching problem for regular expressions extended with any of the four (for backreference, even when restricted to one capturing group) cannot be solved in $O(n^{2-\varepsilon}\operatorname{poly}(m))$ time for any constant $\varepsilon>0$ under the Orthogonal Vectors Conjecture. Moreover, we study the matching problem for regular expressions extended with complement in more detail, which is also known as extended regular expression (ERE) matching. We show that there is no ERE matching algorithm that runs in $O(n^{\omega-\varepsilon}\operatorname{poly}(m))$ time ($2\leq\omega<2.3714$ is the exponent of square matrix multiplication) for any constant $\varepsilon>0$ under the $k$-Clique Hypothesis, and there is no combinatorial ERE matching algorithm that runs in $O(n^{3-\varepsilon}\operatorname{poly}(m))$ time for any constant $\varepsilon>0$ under the Combinatorial $k$-Clique Hypothesis. This shows that the $O(n^{3}m)$-time algorithm introduced by Hopcroft and Ullman in 1979 and recently improved by Bille et al. to run in $O(n^{\omega}m)$ time using fast matrix multiplication was already optimal in a sense, and sheds light on why the theoretical computer science community has struggled to improve the time complexity of ERE matching with respect to $n$ and $m$ for more than 45 years.
@@ 1 Introduction
On the other hand, regular expression matching becomes much harder when extensions are present. A prominent example is backreference, an extension that allows reusing previously matched substrings. The extension is practically popular and supported in the standard libraries of many popular programming languages including Java, Python, JavaScript and more. Unfortunately, we cannot solve the matching problem for regular expressions extended with backreference (rewbs for short) using Thompson’s algorithm because rewbs are strictly more expressive than regular languages (in fact, they are not even context-free) [11] and therefore an equivalent NFA does not in general exist. Instead, it can be solved by simulating memory automata, which are a model equivalent to rewbs introduced by Schmid [51]. Nevertheless, the resulting matching algorithm for rewbs with at most $\ell$ capturing groups takes $O(n^{2\ell+1}m)$ time [51, 52, 45], and at present, even when $\ell=1$, we incur an additional $n^{2}$ factor compared to the running time of Thompson’s algorithm. Another practically popular extension is bounded repetition $r\{i,j\}$, which denotes repeated concatenation of expression $r$ by specifying a lower bound $i$ and an upper bound $j$ on the number of repetitions. In the special case $i=j=2$, the extension is called squaring [40]. Although bounded repetition can be eliminated by unrolling the repetition bounds, the resulting expression can become exponentially larger. Therefore, the naive algorithm eliminating it and running Thompson’s algorithm takes at least exponential time in $m$.
@@
Other well-known extensions that make the matching problem difficult are intersection and complement, where regular expressions extended with each of them are called semi-extended regular expressions (semi-EREs) and extended regular expressions (EREs), respectively. We note that EREs essentially subsume semi-EREs because intersection can be expressed using complement and alternation with only a linear increase in the length of the expression. The best known time complexity of both semi-ERE matching and ERE matching with respect to $n$ and $m$ is $O(n^{\omega}m)$ [15], where $2\leq\omega<2.3714$ [7] is the exponent of square matrix multiplication.
@@
In light of the above, we pose the following question: Do these extensions actually increase the time complexity of regular expression matching, or do they only appear hard because we have not yet discovered the right algorithmic ideas? Notably, lookaround is a practically popular extension to regular expressions like backreference and bounded repetition. The extension can be used to test surrounding contexts of a position and is often used to mimic intersection and complement. Recent studies have shown a surprising fact that this extension does not actually increase the complexity of regular expression matching by presenting $O(nm)$-time algorithms for the matching problem with the extension [39, 24, 10].
1.1 Result I: Hardness of Matching with Extensions under OVC
@@
In this paper, we consider the four extensions described above, backreference, squaring, intersection and complement, and show that regular expression matching extended with each of the four (for backreference, even when restricted to one capturing group) is hard under the Orthogonal Vectors Conjecture (OVC). We note that this also holds under the Strong Exponential Time Hypothesis (SETH) [33] because SETH implies OVC [56, 58].
@@
Unless OVC fails, there is no algorithm that solves rewb matching in $O(n^{2-\varepsilon}\operatorname{poly}(m))$ time for any constant $\varepsilon>0$, even when restricted to one capturing group. [^1] Here, $\operatorname{poly}(m)$ denotes a constant-degree polynomial in $m$.
We refer to regular expressions extended with squaring as RSQs.
Theorem 2.
Unless OVC fails, there is no algorithm that solves RSQ matching in $O(n^{2-\varepsilon}\operatorname{poly}(m))$ time for any constant $\varepsilon>0$.
@@
Theorems 1, 2 and 3 state a much stronger statement than merely that we should not hope for a matching algorithm for rewbs, RSQs or semi-EREs that runs in $O(nm)$ time. Indeed, it states that if we want to solve rewb matching, RSQ matching or semi-ERE matching in truly subquadratic time in $n$ (i.e., $n^{2-\varepsilon}$ for some $\varepsilon>0$), then we are forced to admit a superpolynomial blowup in $m$ unless OVC fails.
The results are of interest not only to theorists but also to practitioners. They provide evidence that the real-world regular expression engines which guarantee worst-case $O(nm)$-time complexity (e.g., that in the de facto standard regular expression library for Rust [^2] The engine guarantees worst-case $O(nm)$-time complexity, although it supports bounded repetition (and hence also squaring) [1]. This is consistent with Theorem 2 (without refuting OVC) because here $m$ denotes the length of the input regular expression after bounded repetition has been eliminated. See the “Untrusted input” section in [1].) are all forced to give up the full support for backreference, squaring and intersection, and would need to somehow restrict their usage such that our reductions in the proofs of Theorems 1, 2 and 3 cannot apply. Thus, the results are of great practical value because they provide guidelines on how far the expressiveness can be increased in regular expression engines while retaining the good complexity.
@@
1.2 Result II: Hardness of Matching with Complement under $k$-Clique Hypotheses
@@ 1.3 Technical Overview
All lower bounds shown in this paper are of the form $O(n^{c-\varepsilon}\operatorname{poly}(m))$ and therefore do not rule out algorithms that run in superpolynomial time in $m$ (e.g., $O(n^{c-1}2^{m})$). Still, this polynomial-in-$m$ lower bound rules out algorithms that are efficient with respect to both $n$ and $m$. Moreover, restricting the time complexity of RSQ matching and ERE matching with respect to $m$ is even necessary for deriving a meaningful lower bound on the complexity with respect to $n$ and $m$ because we could improve the complexity to be linear in $n$ if we were to allow the complexity to be unbounded in $m$. Indeed, the naive RSQ matching algorithm mentioned earlier that eliminates squares by unrolling and uses Thompson’s algorithm runs in time linear in $n$ but exponential in $m$. ERE matching can also be solved in linear time in $n$ by converting a given ERE to an equivalent NFA and running Thompson’s algorithm. But this makes the algorithm’s running time nonelementary in $m$ [53]. Rosu has proposed a less extreme ERE matching algorithm that runs in $O(n^{2}\,2^{m})$ time [49].
We prove our theorems by reductions from the respective problems. A key common technique in our reductions is to exploit the polynomial-in-$m$ lower-bound property to allow the regular expressions to be of polylogarithmic lengths with respect to the sizes of the instances to be reduced. For example, in the proofs of Theorems 1, 2 and 3, we show reductions from the Orthogonal Vectors problem, which asks whether given sets $A$ and $B$ containing the same number of bit-vectors of dimension $d$ have a pair of orthogonal vectors $a\in A$ and $b\in B$. We encode the vectors in $A$ and $B$ into a string and construct a regular expression $r$ that nondeterministically guesses two vectors $a\in A$ and $b\in B$ from the string and checks their orthogonality. In proving a lower bound under OVC, we may assume that $d$ is logarithmic in the number of vectors in $A$ and $B$. Therefore, we can let the length of $r$ be polynomial in $d$ for proving our polynomial-in-$m$ lower bounds. As we shall show, such $r$ can be constructed by exploiting this property and the respective regular expression extensions. A similar argument applies to Theorem 4 where we show a reduction from the $3k$-Clique problem.
@@ 2 Related Work
In this section, to clarify the relation to our work, we express running times from the prior work with respect to the regular expression length $m$ (and the input string length $n$) even when the original bounds were given using more refined parameters of regular expressions, unless stated otherwise. We first discuss the prior work on regular expression matching without extensions. A line of work has shown that the complexity of Thompson’s algorithm can be improved [41, 12, 13, 16, 17, 14]. On the other hand, Backurs and Indyk showed that the problem cannot be solved in $O((nm)^{1-\varepsilon})$ time for any $\varepsilon>0$ unless SETH fails [9]. Later, Abboud and Bringmann introduced a hypothesis, called the Formula-SAT Hypothesis, and showed a more fine-grained lower bound assuming it [3].
@@
Next, we discuss the prior work on rewb matching. Aho showed that the problem is NP-complete [6]. Moreover, it is known that rewbs can be regarded as a generalization of Angluin’s pattern languages (also known as patterns with variables), whose matching problem is already NP-complete [8, 21], even under bounded parameterizations [50, 22]. At this point, one may wonder about the novelty of Theorem 1 because NP-hardness already rules out rewb matching algorithms that run in $O(n^{2-\varepsilon}\operatorname{poly}(m))$ time (in fact, even those that run in time polynomial in $n+m$) assuming $\mathrm{P}\neq\mathrm{NP}$. However, it cannot rule out matching algorithms for rewbs with at most $\ell$ capturing groups for a fixed $\ell$, because the problem is solvable in $O(n^{2\ell+1}m)$ time as stated in the introduction and hence is in $\mathrm{P}$ when $\ell$ is fixed. By contrast, our Theorem 1 can rule out those for rewbs with even one capturing group under OVC. Furthermore, to our knowledge, there is no known implication between OVC and $\mathrm{P}\neq\mathrm{NP}$ (although both follow from SETH [58, 20]) and therefore these hardness results appear to be independent. Regarding upper bounds for the problem, Freydenberger and Schmid introduced a deterministic variant of rewbs together with algorithms that decide whether a given rewb is deterministic in their sense in $O(m^{2})$ time and solve the matching problem for those rewbs in $O(m^{2}+n\ell)$ time [23]. Therefore, combining this with our Theorem 1, we obtain evidence under OVC that their determinism makes rewb matching strictly more tractable (even when $\ell=1$). Terauchi [54] has shown that the matching problem for rewbs satisfying a condition called constant-degree ambiguity (CDA) can be solved in linear time in $n$. They leave open the decidability of checking whether a given rewb satisfies CDA. Additionally, Nogami and Terauchi [45] recently proposed an $O(n^{2}m^{2})$-time algorithm for a certain subclass of rewbs with one capturing group, improving the previously known best $O(n^{3}m)$ algorithm for the subclass with respect to $n$. We comment on the expressive power of rewbs. As mentioned in the introduction, backreference differs from bounded repetition, intersection and complement in that it strictly increases the expressive power of regular expressions. The expressive power of rewbs is known to be incomparable with that of context-free languages (CFLs) [18, 11]. Furthermore, recent papers have shown that it is properly contained in that of EDT0L languages and incomparable with both the class of stack languages and that of multiple context-free languages (MCFLs), while a syntactic condition of rewbs, called the closed-star condition, decreases its expressive power to be properly contained in subclasses of those, namely the class of nonerasing stack languages and that of unary-MCFLs [43, 44, 46].
Next, we discuss the prior work on the matching problem extended with bounded repetition. Kilpeläinen and Tuhkanen showed that the problem is in P [35]. The problem for subclasses of bounded repetition has been extensively studied. Indeed, various formalizations of determinism have been introduced in the literature [36, 30, 34, 31, 25, 19, 27]. One such formalization is strong determinism. Whether a given regular expression with bounded repetition is strongly deterministic can be decided in linear time [19, 27], and the matching problem for strongly-deterministic expressions can be solved in $O(n\log{\log{m}})$ time after a preprocessing phase that runs in $O(m)$ expected time [27]. Glaunec et al. [26] proposed a matching algorithm for a certain subclass with no nested bounded repetition that runs in $O(nm)$ time and with the running time independent of the repetition bounds. Therefore, as in the case of deterministic rewb matching, these prior results and our Theorem 2 provide evidence under OVC that the matching problems for the respective subclasses considered by the prior works are strictly more tractable. [^4] For strongly-deterministic expressions, we strengthen OVC so that it rules out even randomized algorithms. Such a variant can be found in the literature [59]. Also, [27] notes that their preprocessing can be made deterministic. For RSQs, Meyer and Stockmeyer showed that deciding universality (and hence also equivalence) requires exponential space and therefore it is not in P [40]. It is also known that RSQ matching extended with intersection is in LOGCFL [47] and that extended with complement is in P [53].
@@
Next, we discuss the prior work on ERE matching. As mentioned in the introduction, Hopcroft and Ullman [29] proposed in 1979 an algorithm that runs in $O(n^{3}m)$ time and no polynomial improvement over this with respect to $n$ and $m$ had existed for many years prior to Bille et al. [15]. ERE matching is extensively studied and has an interesting history. In fact, many papers had claimed to propose an improved $O(n^{2}m)$-time algorithm [28, 60, 38, 32] ([60] considers semi-ERE matching). However, all of these claims were later found to be incorrect; see [48, 49] for more details. Additionally, a number of studies had improved the complexity of ERE matching with respect to more refined parameters than just $n$ and $m$ [61, 62, 15]. As also mentioned in the introduction, Bille et al. [15] recently adopted fast matrix multiplication to improve the time and space complexity of Hopcroft and Ullman’s algorithm and that of the algorithm by Yamamoto and Miyazaki [62]. Furthermore, as mentioned in Section 1.3, Rosu proposed an algorithm that runs in time quadratic in $n$, but exponential in $m$ [49]. On the hardness side, Petersen proved that semi-ERE matching is LOGCFL-complete by showing a reduction $f$ from every context-free language (CFL) to semi-ERE matching [47]. Combining $f$, which is combinatorial, and the combinatorial reduction from $3k$-Clique to CFL recognition shown by Abboud et al. [2], we can derive the following hardness of semi-ERE matching: (i) there is no semi-ERE matching algorithm that runs in $O((n+m)^{\omega-\varepsilon})$ time for any $\varepsilon>0$ under the $k$-Clique Hypothesis, and (ii) there is no combinatorial semi-ERE matching algorithm that runs in $O((n+m)^{3-\varepsilon})$ time for any $\varepsilon>0$ under the Combinatorial $k$-Clique Hypothesis. However, note that these bounds cannot establish the optimality of Hopcroft and Ullman’s algorithm in the sense described in Section 1.2. By contrast, our Theorem 4 can do it. This is because our reduction outputs EREs of length only polylogarithmic in the input size, whereas $f$ outputs EREs of length linear in the input size.
@@
Finally, we note that our form of multivariate lower bounds is not new. For example, Abboud et al. [5] proved that under OVC, for any $\varepsilon>0$, there is no $n^{2-\varepsilon}\,2^{o(k)}$-time algorithm that decides if the diameter of a given graph of treewidth $k$ is no more than $2$.